OpenStudy (anonymous):

If a 4x4 Matrix A with rows v1, v2, v3 and v4 has a det A = 3, then det [ 6v1+3v4; v2 ; v3; 4v1+8v4] = ???

OpenStudy (zarkon):

\[A=\left( \begin{array}{c} v_1 \\ v_2 \\ v_3 \\ v_4 \\ \end{array} \right)\] with Det(A)=3 \[2r_1\rightarrow r_1\] \[A_1=\left( \begin{array}{c} 2v_1 \\ v_2 \\ v_3 \\ v_4 \\ \end{array} \right)\] \[r_4+r_1\rightarrow r_1\] \[A_2=\left( \begin{array}{c} 2v_1+v_4 \\ v_2 \\ v_3 \\ v_4 \\ \end{array} \right)\] \[6r_4\rightarrow r_4\] \[A_3=\left( \begin{array}{c} 2v_1+v_4 \\ v_2 \\ v_3 \\ 6v_4 \\ \end{array} \right)\] \[2r_1+r_4\rightarrow r_4\] \[A_4=\left( \begin{array}{c} 2v_1+v_4 \\ v_2 \\ v_3 \\ 4v_1+8v_4 \\ \end{array} \right)\] \[3r_1\rightarrow r_1\] \[A_5=\left( \begin{array}{c} 6v_1+3v_4 \\ v_2 \\ v_3 \\ 4v_1+8v_4 \\ \end{array} \right)\] thus the \[Det(A_5)=3\cdot2\cdot6\cdot3=108\] the 2,6,3 come from multiplying a row by a constant.